OpenStudy (anonymous):

Let T : C∞(R) −→ C∞(R) be the map given by T(f) = f′′− 2f′− 3f . Find a linear combination a sin(x) + b cos(x) which solves the equation f′′− 2f′− 3f = 2 cos(x).

OpenStudy (anonymous):

This combination will work \[f(x)=\frac{1}{5} (-\sin (x)-2 \cos (x)) \] To show that you have to say \[ f(x) = A \cos(x) + B \sin(x) \] and find A and B such that f′(x)′− 2f′(x) − 3 f(x) = 2 cos(x).\\ Now \[ f'(x)=B \cos (x)-A \sin (x)\\ f''(x)=-A \cos (x)-B \sin (x)\\ f''(x)− 2f′(x) − 3 f(x)=-2 ((2 A+B) \cos (x)-(A-2 B) \sin (x))= 2\cos(x)\\ \] Equating the coefficient of cos(x) and sin(x), you get \[ 2 A+B =-1 \\ A- 2 B =0 \] You solve for A and B \[A=-\frac{2}{5} \\ B=-\frac{1}{5} \] and you replace.